OpenStudy (acacia):

What does the equation x^2+y^2=14 represent

OpenStudy (acacia):

A. Parabola B. Circle C. Hyperbola D. Ellipse

OpenStudy (anonymous):

\[x^2+y^2=r^2\] This equation represents a circle centered at the origin with radius r. So, your example is a circle centered at the origin with a radius of sqrt{14}

OpenStudy (acacia):

what does the equation \[\frac{ x^2 }{ 81}+\frac{ y^2 }{ 1 }=1\] represent? @vinnv226

OpenStudy (anonymous):

This is similar to a circle, but the 81 under the x^2 creates an ellipse.

OpenStudy (acacia):

would \[\frac{ y^2 }{ 16 }-\frac{ x^2 }{ 9 }=1\] be an hyperbola?

OpenStudy (acacia):

@vinnv226

OpenStudy (anonymous):

Yes, the subtraction sign makes it a hyperbola. If it was a plus sign it would be an ellipse.